# Ask, Reason, Assist: Decentralized Robot Collaboration via Language and Logic

### v1

@@ III-B Translating Help Proposals to STL Specifications
To translate help proposals from NL to STL, We draw from a body of recent work on the transformation of natural language tasks and specifications into temporal logic formulas using large language models [9, 10, 11]. We enhance these natural language to temporal logic methods by defining a Backus-Naur form (BNF) grammar, a notation system for defining formal languages that can be used to constrain the output of LLMs [20]. Additionally, we finetune the LLM model using LoRA [21]. Specifically, defining a BNF grammar for STL involves defining unary and binary temporal and boolean operators, allowed predicates, and text representations of temporal logic relations. BNF constrained generation enforces the syntactic validity of generated temporal logic formulas, allowing us to directly feed LLM outputs into an STL solver [^3] Details on how we encode STL specifications can found in the appendix.. However, constrained generation has been shown to degrade task performance in many cases, i.e. not “minimally invasive” [22]. As such, we carefully designed the BNF grammar [^4] An example BNF grammar can be found in Sec. -A3. to be sufficiently relaxed as to be “minimally invasive”, for example including optional whitespace and parentheses, and allowing for arbitrary nesting and recursion, while maintaining the guarantee that the output can be processed using standard STL parsers. Ultimately, we generate an STL specification by reasoning over the help offer message and conflict location, conditioned on a prompt $p^{\mathrm{\textnormal{STL}}}$ and subject to the BNF grammar:
@@ III-C Solving for Optimal Robot Paths
Suppose an initial, globally near-optimal task schedule $S^{\mathrm{\textnormal{orig}}}=\{\Psi_{0},...,\Psi_{n-1}\}$ has been determined. Each candidate helper robot $j\in\mathcal{R}_{-i}$ has an original task specification $\varphi^{\mathrm{\textnormal{orig}}}_{j}$ and has found a corresponding optimal path
@@
by minimizing Manhattan distance [^5] We use Manhattan distance to capture the number of movements in a grid world and improve computation time over solving a Mixed Integer Quadratic Program (MIQP) with Euclidean Distances. and time to satisfy the task specification (i.e., makespan) while always obeying global specifications:
@@
Once the help offer $m^{\mathrm{\textnormal{o}}}_{j\to i}$ is translated into $\varphi^{\mathrm{\textnormal{h}}}_{j}$ as in 8, robot $j$ computes an updated path that both helps and completes the original tasks:
@@
We formulate 10 and 11 as Mixed Integer Linear Programs (MILPs) by defining predicates in the STL formula as binary variables over discrete time steps, and solve the formulated MILPs with Gurobi [18].
@@
Note that, in our implementation, $\tau^{\mathrm{\textnormal{h}}}_{j}$ and $\tau^{\mathrm{\textnormal{new}}}_{j}$ are computed as a byproduct of solving 11, instead of solving new MILPs to compute these makespans.
Importantly, our method does not alter the original task schedule $S^{\mathrm{\textnormal{orig}}}$; only the chosen helper $j$ augments its plan with the help task. This contrasts with an approach using a centralized Oracle (c.f., [15, 16]), which is free to re-allocate any task from any $\psi_{k}$ to any other robot to minimize the total system cost which is the sum of the individual makespan, $\sum_{i\in\mathcal{R}}M(x_{i}(0),\varphi_{i})$. This means that our method sacrifices optimality for decentralization, as we assess next.
@@ IV Experiments
To demonstrate the efficacy of the proposed end-to-end framework for multi-robot collaboration, we isolate and evaluate its two constituent modules: the natural language–to–temporal logic translation component and the temporal logic–to–task plan generation component. Each module is quantitatively compared against state-of-the-art methods in natural language–to–temporal logic translation and vehicle routing, respectively.
All experiments were performed using Python [^6] Code and Fine-tuned LLM model available at:
https://github.com/ask-reason-assist/ask-reason-assist-codebase
on a desktop computer with a 32-core i9 CPU, 32 GB RAM, and an NVIDIA RTX 4090 GPU.
@@ IV-A Experiment 1: Natural Language to Temporal Logic
To benchmark NL to TL translation performance, we use the dataset presented in [9], which consists of 7,500 pairs of natural language sentences and corresponding signal temporal logic (STL) formulas in the context of navigation tasks. 6,000 pairs are used for finetuning, and 1,500 pairs are used for evaluation.
We compare our method to several ablated variants and to a GPT-4 baseline presented in [9]. The ablations we consider include variants without the grammar constraint, without the grammar prompt, and with combinations thereof, allowing us to isolate the contributions of each component of our method. Our method and each ablation are evaluated with 5 and 20 “few-shot” examples of natural language tasks with accurate temporal logic translations included in the prompt to the LLM, with 500 NL-TL pairs randomly sampled from the dataset. Each evaluation is rerun 3 times, with resampled few-shot examples and NL-TL pairs. In our approach, both the BNF grammar and few shot examples are included in the prompt, and the BNF grammar is enforced during decoding using llama.cpp’s constrained generation. Results with constrained generation are not presented for GPT-4, as this model can only be accessed through the OpenAI API, which does not allow for a BNF grammar constraint to be specified.
@@
We measure Validity as the percentage of generated formulas that are syntactically correct and Accuracy as percentage of valid generated STL formulas logically equivalent to true STL formulas from the dataset. Accuracy is computed on the mutually valid set of generated formulas across all ablations. The logical equivalence and syntactic correctness of STL formulas are checked using the Spot library [24]. Note we also measure Inference Time as average runtime per NL-STL pair (ms), in Sec. -C.
@@
The results are summarized in Table I. Our method achieves 100% formula validity in all cases with no loss in accuracy due to constrained generation, confirming our hypothesized guarantee. Additionally, we show improved translation accuracy in the navigation dataset with a significantly smaller LLM.
@@
See Table III for the inference time of our method and each ablation.
@@
5 Gemma F + P + C (Ours) 100.0 $\pm$ 0.00 98.98 $\pm$ 0.17
  Gemma F + P 99.8 $\pm$ 0.45 98.98 $\pm$ 0.17
  Gemma F 99.0 $\pm$ 1.22 84.97 $\pm$ 8.65
GPT-4 F + P 100.0 $\pm$ 0.00 62.82 $\pm$ 3.66
20 Gemma F + P + C (Ours) 100.0 $\pm$ 0.00 97.58 $\pm$ 1.16
  Gemma F + P 99.67 $\pm$ 0.01 97.58 $\pm$ 1.16
  Gemma F 99.60 $\pm$ 0.04 93.73 $\pm$ 2.40
GPT-4 F + P 100.0 $\pm$ 0.00 92.73 $\pm$ 3.00
@@
5 Gemma F + P + C (Ours) 99.45 $\pm$ 0.24
  Gemma F + P 99.45 $\pm$ 0.24
  Gemma F 87.02 $\pm$ 9.19
GPT-4 F + P 74.13 $\pm$ 6.62
20 Gemma F + P + C (Ours) 99.25 $\pm$ 0.23
  Gemma F + P 99.25 $\pm$ 0.23
  Gemma F 97.52 $\pm$ 0.89
GPT-4 F + P 95.20 $\pm$ 6.64
TABLE II: Containment checking of LLM generated STL formulas for each NL to TL ablation. If the true STL formula is contained in the LLM-generated STL formula, the constraint specified in the natural language will always be satisfied.
@@ IV-B Experiment 2: Mobile Robot Blocked by a Pallet
In our second experiment, we compare our decentralized framework to a centralized “Oracle” in a scenario where forklift robots respond to a help request from a mobile robot prompted by the scene description $m^{\mathrm{\textnormal{c}}}=$ “A pallet is blocking the entrance to the picking aisle.” Each forklift robot $j\in\mathcal{H}$ then evaluates whether it can handle this additional task (moving the obstructing pallet) on top of its existing pick-and-place (PNP) [^7] Our STL specification of a PNP job is in Sec. -A2. jobs by updating its STL specification $\varphi^{\mathrm{\textnormal{orig}}}_{j}$. The help task, $\varphi_{j}^{\mathrm{\textnormal{h}}}$, is similarly encoded as a PNP job where the forklift must travel to the help site, pick up the obstructing pallet, and place the pallet in the nearest free cell $x^{\mathrm{\textnormal{adj}}}=\arg\min_{x}\{\|x-x^{c}\|_{1}\}$.
@@
We conducted 100 simulation trials with randomly chosen help-site locations. To ensure a fair comparison within each trial, a single, consistent starting scenario was established for the fleet.
This involved spawning six forklift robots at random initial positions and distributing total of twelve PNP tasks among the fleet according to a global schedule, $S^{\mathrm{\textnormal{orig}}}$. We pre-compute the schedule using the “Oracle” baseline to find a near-optimal solution to the MRTA problem. When a help task $\varphi^{\mathrm{\textnormal{h}}}$ arises, the Oracle finds a new via ILS metaheuristics, as detailed [^8] The algorithm table can be found in Sec. -D in [16].
Our method and the three distinct baselines were then evaluated against this identical initial setup: (B1) The ILS-based global planner “Oracle”, (B2) the closet forklift to the help site solves for the optimal path via MILP, and (B3) a hybrid approach where help task is assigned to the closet agent, but the Oracle is allowed to re-optimize all other tasks.
@@
Fig. 2 summarizes our findings. The Centralized Oracle (B1) added an average of 4.49 time-steps to the system by inserting the help task, establishing a strong performance benchmark. Our decentralized method achieved a highly competitive result of 5.47 steps on average with a standard deviation of 3.32 time steps, supporting that informed local optimization captures the majority of the performance gains of a centralized system without the need for global task reassignment or heavy information demand. Importantly, the average Gurobi model solve time was 5.3 seconds with 30 discretized time horizon, demonstrating computational practicality for real-time deployment.
On the other hand, our optimal cost approach on average provides 46% and 53% efficiency gains over B2 and B3 respectively. We observe that the inter-dependencies in PNP jobs make distance-based heuristics (B2-B3) suboptimal, consistent with previous literature in operations research [25]. Furthermore, an initial myopic decision (B3) can constrain the overall system’s optimality from which even a powerful optimizer cannot recover efficiently [26]. In contrast, our method explicitly avoids such myopia by reasoning over future task interactions.
Figure 2: Comparison of total time-steps added to the system under different methods tested in Sec. IV-B. Our decentralized method tracks the Oracle solution within 18% while significantly outperforming the distance based heuristics (B2) and hybrid approach (B3).
@@ V Demonstrations
We now demonstrate the utility of NL in conflict resolution on three complex warehouse tasks. These complement our experiments, which stress-test components of our method. Here, we illustrate the practicality of our overall framework by implementing it in a high-fidelity physics-based Unity simulator. In particular, we integrate our NL-to-TL translation module with a MILP-based TAMP mechanism for executing sequentially nested tasks, alongside local collision avoidance and an A* path planner for realistic motion plans. Full videos of each demonstration are available on our project website https://ask-reason-assist.github.io.
@@
Our second scenario is a warehouse kitting task, a common workflow in fulfillment operations. Here the NL input is, “Eventually pick up an item from Aisle A and Aisle B and Aisle C.” Our NL-TL module successfully translates this into a semantically correct and syntactically valid TL formula:
@@
In our final, most complex scenario, the NL help request is, “First visit the front office to pick up the high fidelity scanner, then travel to the help site to scan the damaged good, and finally navigate back to the front office to return the scanner.” This is translated into the nested STL formula:
@@ VI Conclusion and Future Work
Each robot is subject to global safety and actuation constraint. For example, all ground vehicles must avoid obstacle cells and can only move one grid space in any cardinal direction):
@@
-C Experiment I Inference Time Results
# Examples Method Variant Inference Time (ms)
5 Gemma F + P + C (Ours) 801.6 $\pm$ 36.2
  Gemma F + P 583.4 $\pm$ 28.2
  Gemma F 410.2 $\pm$ 9.2
GPT-4 F + P 803.7 $\pm$ 30.3
20 Gemma F + P + C (Ours) 1191.7 $\pm$ 14.1
  Gemma F + P 974.9 $\pm$ 2.5
  Gemma F 414.9 $\pm$ 3.9
GPT-4 F + P 2736.9 $\pm$ 25.2
TABLE III: Inference times for NL-to-TL method variants with varying numbers of few-shot examples.
-D Oracle Baseline with the ILS Algorithm

### v3

@@ III-B Translating Help Proposals to STL Specifications
To translate help proposals from NL to STL, we draw from recent work using LLMs [5, 12, 19]. We enhance these NL-to-TL methods by defining a Backus-Naur form (BNF) grammar, a notation system for defining formal languages that can be used to constrain the output of LLMs [26]. Additionally, we finetune the LLM model using LoRA [15]. Specifically, defining a BNF grammar for STL involves defining unary and binary temporal and boolean operators, allowed predicates, and text representations of temporal logic relations. BNF constrained generation enforces the syntactic validity of generated temporal logic formulas, allowing us to directly feed LLM outputs into an STL solver (details on how we encode STL specifications can found in the appendix). However, constrained generation has been shown to degrade task performance when grammar constraints are misaligned with the model’s tokenization [3]. As such, we carefully design the BNF grammar (example in App.-A3) to be sufficiently relaxed, for example including optional whitespace and parentheses, and allowing for arbitrary nesting and recursion, while maintaining the guarantee that the output can be processed using standard STL parsers. Ultimately, we generate an STL specification by reasoning over the help offer message and conflict location, conditioned on a prompt $p^{\mathrm{\textnormal{STL}}}$ and subject to the BNF grammar:
@@
Note, enforcing BNF grammar with constrained generation guarantees syntactically valid temporal logic formulas, but not semantic equivalence to the natural language specification. We evaluate the impact of this limitation in Sec. IV-A.
@@ III-C Solving for Optimal Robot Paths
Suppose an initial task schedule $S^{\mathrm{\textnormal{orig}}}=\{\Psi_{0},...,\Psi_{n-1}\}$ that approximately minimizes total system makespan has been determined. Each candidate helper $j\in\mathcal{R}_{-i}$ has an original specification $\varphi^{\mathrm{\textnormal{orig}}}_{j}$ and corresponding optimal path
@@
by minimizing Manhattan distance (simplifies computation over Euclidean distance) and time to satisfy the task specification (i.e., makespan) while always obeying global specifications:
@@
Once the help offer $m^{\mathrm{\textnormal{o}}}_{j\to i}$ is translated into $\varphi^{\mathrm{\textnormal{h}}}_{j}$ as in 6, robot $j$ computes an updated path that enables it to both help and complete the original tasks:
@@
We formulate 8 and 9 as MILPs by defining predicates in the STL formula as binary variables over discrete time steps, and solve with Gurobi [14] which provides the spatio-temporal grounding that NL help offer $m^{\mathrm{\textnormal{o}}}_{j\to i}$ lacks.
@@
Note that, in our implementation, $\tau^{\mathrm{\textnormal{h}}}_{j}$ and $\tau^{\mathrm{\textnormal{new}}}_{j}$ are computed as a byproduct of solving 9, instead of solving new MILPs to compute these makespans.
Importantly, our method does not alter the original task schedule $S^{\mathrm{\textnormal{orig}}}$; only the chosen helper $j$ augments its plan with the help task. This contrasts with an approach using a centralized Oracle (cf., [21, 18]), which is free to re-allocate any task from any $\psi_{k}$ to any other robot to minimize the total system cost, defined as the sum of the individual makespan, $\sum_{i\in\mathcal{R}}M(x_{i}(0),\varphi_{i})$. This means that our method sacrifices optimality for scalability and minimal information disclosure, as we assess next.
@@ IV Experiments
To study the efficacy of the proposed end-to-end framework for multi-robot collaboration, we isolate and evaluate its two constituent modules: the natural language–to–temporal logic translation component and the temporal logic–to–task plan generation component. Each module is quantitatively compared against state-of-the-art methods in natural language–to–temporal logic translation and vehicle routing, respectively.
All experiments were performed using Python [^1] Code available online (click for link). on a desktop computer with a 32-core i9 CPU, 32 GB RAM, and an NVIDIA RTX 4090 GPU.
@@ IV-A Experiment 1: Natural Language to Temporal Logic
To benchmark NL to TL translation performance, we use the dataset presented in [5], which consists of 7,500 pairs of natural language sentences and corresponding signal temporal logic (STL) formulas in the context of navigation tasks. 4,500 pairs are used for finetuning, and 3,000 pairs are used for evaluation.
We compare our method to several ablated variants and to a GPT-4 baseline presented in [5]. The ablations we consider include variants without the grammar constraint, without the grammar prompt, and with combinations thereof, allowing us to isolate the contributions of each component of our method. Our method and each ablation are evaluated with 5 and 20 “few-shot” examples of natural language tasks with accurate temporal logic translations included in the prompt to the LLM, with 1000 NL-TL pairs randomly sampled from the dataset. Each evaluation is rerun 3 times, with resampled few-shot examples and NL-TL pairs. In our approach, both the BNF grammar and few shot examples are included in the prompt, and the BNF grammar is enforced during decoding using llama.cpp’s constrained generation. Results with constrained generation are not presented for GPT-4, as this model can only be accessed through the OpenAI API, which does not allow for a BNF grammar constraint to be specified.
@@
We measure Validity as the percentage of generated formulas that are syntactically correct and Accuracy as percentage of valid generated STL formulas logically equivalent to true STL formulas from the dataset. Accuracy is computed on the mutually valid set of generated formulas across all ablations. The logical equivalence and syntactic correctness of STL formulas are checked using the Spot library [9].
@@
The results are summarized in Table I. Our method achieves 100% formula validity in all cases with no loss in accuracy due to constrained generation, confirming our hypothesized guarantee. Constrained generation results in a modest increase in inference time of 218.2 ms on average. Additionally, we show improved translation accuracy in the navigation dataset with a significantly smaller LLM.
While semantic accuracy generally improves with advancements in large language models, adhering to the rigid, domain-specific formal language syntax remains a distinct problem. Additionally, although large models like GPT-4, which has approximately 1.8 trillion parameters compared to our model’s 12 billion parameters, can often produce valid syntax without explicit grammar constraints, their computational requirements prevent local deployment. The proposed method ensures syntactic correctness independent of model capacity, facilitating the use of smaller models for edge deployment.
@@
5 Gemma F + P + C (Ours) 100.0 $\pm$ 0.00 99.24 $\pm$ 0.62
  Gemma F + P 99.53 $\pm$ 0.31 99.11 $\pm$ 0.82
  Gemma F 93.73 $\pm$ 6.41 92.94 $\pm$ 3.88
GPT-4 F + P 99.87 $\pm$ 0.12 63.83 $\pm$ 8.15
20 Gemma F + P + C (Ours) 100.0 $\pm$ 0.00 98.44 $\pm$ 0.1
  Gemma F + P 99.40 $\pm$ 0.53 98.03 $\pm$ 0.26
  Gemma F 99.20 $\pm$ 0.40 93.54 $\pm$ 1.97
GPT-4 F + P 99.87 $\pm$ 0.23 93.81 $\pm$ 2.04
@@
5 Gemma F + P + C (Ours) 99.73 $\pm$ 0.47
  Gemma F + P 99.73 $\pm$ 0.31
  Gemma F 97.64 $\pm$ 1.68
GPT-4 F + P 75.15 $\pm$ 12.15
20 Gemma F + P + C (Ours) 99.05 $\pm$ 0.43
  Gemma F + P 98.77 $\pm$ 0.74
  Gemma F 96.12 $\pm$ 2.01
GPT-4 F + P 96.46 $\pm$ 2.33
TABLE II: Containment checking of LLM generated STL formulas for each NL to TL ablation. If the true STL formula is contained in the LLM-generated STL formula, the constraint from natural language will always be satisfied.
@@ IV-B Experiment 2: Mobile Robot Blocked by a Pallet
In our second experiment, we compare our decentralized framework to a centralized “Oracle” in a scenario where forklift robots respond to a help request from a mobile robot prompted by the scene description $m^{\mathrm{\textnormal{c}}}=$ “A pallet is blocking the entrance to the picking aisle.” Each forklift robot $j\in\mathcal{H}$ then evaluates whether it can handle this additional task (moving the obstructing pallet) on top of its existing pick-and-place (PNP) [^2] Our STL specification of a PNP job is in App.-A2. jobs by updating its STL specification $\varphi^{\mathrm{\textnormal{orig}}}_{j}$. The help task, $\varphi_{j}^{\mathrm{\textnormal{h}}}$, is similarly encoded as a PNP job where the forklift must travel to the help site, pick the obstructing pallet, and place it in the nearest free cell.
@@
We conducted 100 simulation trials with randomly chosen help-site locations, with a single starting scenario to ensure fair comparison within each trial. We spawn six forklift robots at random initial positions and distribute twelve PNP tasks among them per a global schedule, $S^{\mathrm{\textnormal{orig}}}$. We pre-compute the schedule using the “Oracle” baseline to find a near-optimal solution. When a help task $\varphi^{\mathrm{\textnormal{h}}}$ arises, the Oracle finds a new schedule via Iterated Local Search (ILS) metaheuristics, as detailed [^3] The algorithm table can be found in App.-C in [18]. Our method and the three distinct baselines were evaluated against this identical initial setup: (B1) The ILS-based global planner “Oracle”, (B2) the closest forklift to the help site solves for the optimal path via MILP, and (B3) a hybrid approach where the help task is assigned to the closest agent, but the Oracle re-optimizes all other tasks.
@@
Fig. 2 summarizes our findings. The Centralized Oracle (B1) added an average of 4.49 time-steps to the system by inserting the help task, whereas our method achieved a mean of 5.47 time-steps, within 18% of the centralized baseline. This demonstrates that informed local optimization captures most of the performance gains without global task reassignment or full information disclosure. Importantly, the average Gurobi model solve time was 5.3 seconds with 30 discretized time horizon, demonstrating computational practicality for real-time deployment.
On the other hand, our method which selects the helper offering the lowest total cost ($\tau^{\mathrm{\textnormal{h}}}_{j}+\tau^{\mathrm{\textnormal{new}}}_{j}$) provides 46% and 53% efficiency gains over B2 and B3 respectively. We observe that the inter-dependencies in PNP jobs make distance-based heuristics (B2-B3) suboptimal, consistent with previous results in operations research [8]. Furthermore, an initial myopic decision (B3) can constrain the overall system’s optimality from which even a powerful optimizer cannot recover efficiently [6]. In contrast, our method explicitly avoids such myopia by reasoning over future task interactions.
Figure 2: Box plot comparison of total time-steps added to the system under different methods tested in Sec. IV-B. Our method tracks the Oracle solution within 22% (mean) while significantly outperforming the distance based heuristics (B2) and hybrid approach (B3).
@@ V Demonstrations
We now demonstrate the utility of NL in conflict resolution on three complex warehouse tasks. These complement our experiments, which stress-test components of our method. Here, we illustrate our framework’s practicality in a high-fidelity physics-based Unity simulator. In particular, we integrate our NL-to-TL translation module with a MILP-based TAMP mechanism for executing sequentially nested tasks, alongside local collision avoidance and an A* path planner for realistic motion plans. Full videos of each demo are available on our project website.
@@
Our second scenario is a warehouse kitting task, common in fulfillment operations. Here the NL input is, “Eventually pick up an item from Aisle A (Green Kit) and Aisle B (Blue Kit) and Aisle C (Red Kit)”. Our NL-TL module successfully translates this into a semantically correct and syntactically valid TL formula:
@@
In our final, most complex scenario, the request is, “First visit the front office to pick up the high fidelity scanner, then travel to the help site to scan the damaged good, and finally navigate back to the front office to return the scanner,” translated to STL as
@@ VI Conclusion and Future Work
Each robot is subject to global safety and actuation constraints. For example, all ground vehicles must avoid obstacle cells and can only move one grid space in any cardinal direction):
@@
-C Oracle Baseline with the ILS Algorithm